OpenStudy (anonymous):

sqrt u - 5 = 7 medallsssss

OpenStudy (anonymous):

\[\sqrt{u-5}=7\]

OpenStudy (anonymous):

thats the question ^

OpenStudy (anonymous):

What number minus 5 equals 7

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

\[x - 5 =7\] Can you solve that?

OpenStudy (anonymous):

12?

OpenStudy (anonymous):

Yes, so 12- 5 =7. With that known, you have to find the square root of a number that equals 12. Do you know one?

OpenStudy (anonymous):

\[\sqrt{x}=12\] Can you solve for x?

OpenStudy (anonymous):

is it 6?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

idk then how?

OpenStudy (anonymous):

The square root of what number equals 12?

OpenStudy (anonymous):

\[12 \times 12=\]

OpenStudy (anonymous):

Multiply 12 x 12 :)

OpenStudy (anonymous):

144

OpenStudy (anonymous):

144 is the answer?

OpenStudy (anonymous):

Yes, you are correct. 144 is the answer your looking for. :)

OpenStudy (anonymous):

144 is the answer to the problem?

OpenStudy (anonymous):

Yes.